### v1

# Robust Cross-View Geo-Localization via Content-Viewpoint Disentanglement

Abstract
Cross-view geo-localization (CVGL) aims to match images of the same geographic location captured from different perspectives, such as drones and satellites. Despite recent advances, CVGL remains highly challenging due to significant appearance changes and spatial distortions caused by viewpoint variations. Existing methods typically assume that cross-view images can be directly aligned within a shared feature space by maximizing feature similarity through contrastive learning. Nonetheless, this assumption overlooks the inherent conflicts induced by viewpoint discrepancies, resulting in extracted features containing inconsistent information that hinders precise localization. In this study, we take a manifold learning perspective and model the feature space of cross-view images as a composite manifold jointly governed by content and viewpoint information. Building upon this insight, we propose CVD, a new CVGL framework that explicitly disentangles content and viewpoint factors. To promote effective disentanglement, we introduce two constraints: (i) An intra-view independence constraint, which encourages statistical independence between the two factors by minimizing their mutual information. (ii) An inter-view reconstruction constraint that reconstructs each view by cross-combining content and viewpoint from paired images, ensuring factor-specific semantics are preserved. As a plug-and-play module, CVD can be seamlessly integrated into existing geo-localization pipelines. Extensive experiments on four benchmarks, i.e., University-1652, SUES-200, CVUSA, and CVACT, demonstrate that CVD consistently improves both localization accuracy and generalization across multiple baselines.
11footnotetext: Corresponding author: wangdi@xidian.edu.cn
1 Introduction
The widespread deployment of autonomous vehicles, drones, and other intelligent systems has placed growing demands on navigation and localization technologies. To ensure safe and reliable operation, high-precision positioning services have become a fundamental requirement. Cross-view geo-localization (CVGL), an onboard technique independent of external communication infrastructure, offers a promising solution by estimating absolute geospatial coordinates in the absence of conventional localization signals (e.g., GPS) [18, 3, 25, 8]. Given a query image, typically captured from a ground or drone perspective, the goal is to find a matching satellite image from a georeferenced database to infer the query’s location. Most existing approaches formulate CVGL as an image retrieval task [17, 23], training deep neural networks (DNNs) to learn visual similarity across different views. However, the viewpoint disparity between query and satellite imagery introduces severe spatial distortions and appearance variations, making robust matching inherently challenging.
Figure 1: Comparison between previous methods and our CVD. Left: Existing methods can be interpreted as operating on a single manifold $\mathcal{M}$, where contrastive objectives directly pull positive pairs closer and push negative pairs away. Right: Our method learns disentangled representations by mapping inputs onto two submanifolds corresponding to content $\mathcal{M}_{c}$ and viewpoint $\mathcal{M}_{v}$. This separation is enforced via two constraints (see Section 3.3 and Section 3.4), promoting effective disentanglement and thereby enhancing cross-view matching performance.
Recent efforts aim to alleviate the viewpoint discrepancy between query and satellite images. A common strategy is to apply predefined geometric transformations to satellite images, such as polar transformation or orthorectification, aligning their spatial layout with that of the query view. However, the effectiveness of these methods relies on prior knowledge of the geometric relationship between views, and may degrade when the query is not spatially centered within the satellite image [31].
Despite recent advances in CVGL, many existing methods [9, 22, 31, 4, 27] still follow a common training paradigm, i.e., directly applying contrastive learning to pull the features of positive pairs closer and push those of negative pairs away. While various architectural or training refinements have been proposed, these methods largely overlook the semantic inconsistencies introduced by drastic viewpoint differences. These inconsistencies disrupt the alignment between positive pairs and ultimately limit the cross-view matching performance.
In this paper, we revisit the CVGL task from a manifold learning perspective by modeling the feature space of cross-view images as a composite manifold jointly governed by content and viewpoint factors. As illustrated in Figure 1 (left), prior methods can be interpreted as learning representations on a composite manifold $\mathcal{M}$, where both factors are inherently entangled. Such entanglement introduces viewpoint-induced conflicts into the learned representations, thereby undermining the robustness of contrastive alignment. To address this limitation, we propose CVD (Content-Viewpoint Disentanglement), a generic CVGL framework that explicitly factorizes the feature space into two submanifolds: a content submanifold and a viewpoint submanifold (see Figure 1 (right)). The content encodes view-agnostic structural information, while the viewpoint captures view-specific appearance variations. As shown in Figure 2 (Training Stage), CVD adopts an embed-disentangle-reconstruct paradigm. Each image is first embedded into a shared feature space, then projected onto independent content and viewpoint submanifolds, and finally recombined in an image reconstruction task. To facilitate effective disentanglement, we impose two dedicated constraints: an intra-view independence constraint and an inter-view reconstruction constraint. The former encourages statistical independence between content and viewpoint by minimizing their mutual information, while the latter ensures that each factor retains its intended semantics by reconstructing images using the content of one view and the viewpoint of the other. In addition, we apply a standard contrastive loss (e.g., InfoNCE) to align content representations of matched pairs across views.
To the best of our knowledge, CVD is the first CVGL framework to explicitly disentangle content and viewpoint information. Unlike existing methods that rely on architectural redesign, CVD serves as a plug-and-play module that can be seamlessly integrated into existing pipelines. It introduces no additional cost during inference while substantially improving matching accuracy and robustness. In summary, our contributions are as follows:
  • We revisit the cross-view geo-localization (CVGL) task from a manifold learning perspective and propose CVD, the first framework that explicitly disentangles content and viewpoint to suppress viewpoint-induced conflicting information.
  • We introduce two constraints to facilitate effective disentanglement: an intra-view independence constraint that facilitates the independence between content and viewpoint, and an inter-view reconstruction constraint that ensures each factor retains its intended semantics.
  • CVD can be seamlessly integrated into existing pipelines, incurs no additional inference overhead, and consistently improves cross-view matching performance and generalization across multiple challenging benchmarks.
2 Related Work
CVGL has witnessed significant progress with the availability of large-scale geo-tagged datasets and advances in deep learning. Existing methods predominantly follow a Siamese-based framework and can be broadly categorized into three research directions: data augmentation strategies, architectural innovations, and feature representation learning.
2.1 Data Augmentation Strategies
Data augmentation has become a widely adopted strategy in CVGL tasks [19, 12, 26, 6]. To address cross-view misalignment, Liu et al.[9] incorporate camera orientation as auxiliary input, while SAFA[15] applies a polar transformation to align aerial and ground-level panoramas. Several methods further improve robustness by diversifying input image appearances. For instance, Lv et al.[10] simulate extreme weather conditions by injecting synthetic variations into University-1652. More recently, training-aware data sampling has emerged as a complementary strategy. Sample4Geo[4] introduces two curriculum-driven strategies: one leveraging geographically adjacent samples for easier early-stage alignment, and another mining hard negatives to refine the decision boundary. Game4Loc [5] proposes a mutual-exclusion sampling mechanism that enforces strict decorrelation between positive and negative pairs, thereby enhancing contrastive supervision in cross-view matching.
2.2 Backbone Innovations
In parallel, a line of work focuses on designing more powerful visual backbones to enhance localization performance [13, 29]. For example, L2LTR [22] exploits self-attention to model long-range dependencies, effectively reducing visual ambiguity in cross-view. RK-Net [7] introduces a lightweight unit-difference attention module that enables joint learning of dense features and salient keypoints, without requiring additional annotations. SAIG [33] proposes an efficient backbone tailored for CVGL by replacing the MLP blocks in standard Transformers with spatially-aware mixing layers and low-dimensional projections, yielding a more compact and structured representation.
2.3 Feature Representation Learning
Many studies focus on learning more effective visual representations to enhance cross-view matching [32, 11, 30, 24]. A common strategy involves refining alignment mechanisms between views. For instance, Shi et al.[16] proposed a dynamic similarity matching network to estimate directional alignment, thereby reducing cross-view discrepancies. FSRA[2] leverages transformer-based heatmaps to perform region-level alignment, while LPN [20] incorporates contextual cues via a square-ring partitioning strategy to improve part-based representations. SDPL [1] builds upon LPN by introducing a shifting-fusion mechanism to generate multiple complementary part sets, which are then adaptively aggregated to enhance robustness against spatial shifts and scale variations. Several methods also incorporate inductive priors modeling to enhance feature expressiveness. TransGeo [31] adopts a non-uniform cropping strategy that discards low-information regions while reallocating resolution to semantically salient areas, enhancing accuracy without increasing computational cost. GeoDTR [27] incorporates spatial layout priors through a geometry-aware extraction module, combined with counterfactual training to exploit structural information. MCCG [14] enriches feature diversity by jointly modeling spatial and channel-wise attentions.
Although prior methods perform well across various scenarios, many rely on auxiliary components that increase inference-time overhead. In contrast, we propose a plug-and-play training paradigm that explicitly disentangles content and viewpoint from raw features, enhancing cross-view correspondence without introducing additional parameters or latency at inference.
Figure 2: Overview of the proposed CVD.
3 Methodology
3.1 Problem Formulation
Considering a set of image pairs $\{(\mathcal{I}_{i}^{\mathrm{q}},\mathcal{I}_{i}^{\mathrm{s}})\}_{i=1}^{N}$, where superscripts $\mathrm{q}$ and $\mathrm{s}$ denote query (e.g., drone or ground) and satellite images, respectively, and $N$ is the number of pairs. Each pair depicts the same geographic location. In the CVGL task, given a query image $\mathcal{I}_{q}^{\mathrm{q}}$ with index $q$, the objective is to retrieve its best-matching satellite image $\mathcal{I}_{b}^{\mathrm{s}}$ from the georeferenced database, where $b\in\{1,\ldots,N\}$.
Most existing methods rely on learning a representation function $f(\cdot)$ that embeds images from different viewpoints into a shared feature space, allowing matching pairs to be identified through feature distance minimization. However, such representations inevitably retain view-specific conflicts, which hinder cross-view semantic alignment and degrade matching performance.
For effective comparison between cross-view images, we aim to modulate raw representations by explicitly disentangling two factors: content and viewpoint. Concretely, we model the feature space $\mathcal{Z}$ as a representation of composite manifold $\mathcal{M}$ structured by two independent submanifolds, $\mathcal{M}_{c}$ and $\mathcal{M}_{v}$, corresponding to content and viewpoint, respectively. Each feature representation is viewed as a sample from two latent random variables, $C\sim p(C)$ and $V\sim p(V)$, defined over $\mathcal{M}_{c}$ and $\mathcal{M}_{v}$. These are composed via a function $f:\mathcal{M}_{c}\times\mathcal{M}_{v}\rightarrow\mathcal{Z}$ that maps both factors to a point in the feature space. Assuming statistical independence between the factors, i.e., $p(C,V)=p(C)\,p(V)$, the resulting distribution over $\mathcal{Z}$ is given by the push-forward measure $f_{\#}(p(C)\times p(V))$. This assumption is well-aligned with the CVGL task, where identical scenes may be observed under diverse perspectives.
To realize the above formulation, we design manifold encoders that decompose each image $\mathcal{I}^{\mathrm{u}}$ into two distinct representations: a content embedding $f_{c}(\mathcal{I}^{\mathrm{u}})\in\mathcal{M}_{c}$ and a viewpoint embedding $f_{v}(\mathcal{I}^{\mathrm{u}})\in\mathcal{M}_{v}$. Cross-view matching is subsequently performed in the $\mathcal{M}_{c}$ by retrieving the nearest neighbor of a query image $\mathcal{I}_{q}^{\mathrm{q}}$ via its content embedding:
$b=\underset{i\in\{1,\ldots,N\}}{\arg\min}d(f_{c}(\mathcal{I}_{q}^{\mathrm{q}}),f_{c}(\mathcal{I}_{i}^{\mathrm{s}})),$ (1)
where $d(\cdot,\cdot)$ denotes a distance metric. For notation compactness, we will use superscript $u$  [^2] We adopt this convention throughout the paper. for cases that apply to both query ($\mathrm{q}$) and satellite ($\mathrm{s}$) views.
3.2 Proposed Methodology
As illustrated in Figure 2, CVD adopts a Siamese architecture consisting of two symmetric branches for the query and satellite views. Each branch comprises three sequential components: manifold embedding, information disentanglement, and cross-reconstruction.
Manifold Embedding. Each input image $\mathcal{I}^{\mathrm{u}}$ ($u\in\{\mathrm{q},\mathrm{s}\}$) is first processed by manifold encoders $E$, yielding a raw $d$-dimensional feature representations $\mathbf{z}^{\mathrm{u}}=E(\mathcal{I}^{\mathrm{u}})\in\mathbb{R}^{d}$. Owing to the nature of DNN encoders, the distribution of $\mathbf{z}^{\mathrm{u}}$ can be viewed as residing on a composite manifold jointly governed by content and viewpoint information. Since our primary focus is on the training paradigm, we adopt the same DNN encoders as those used in the respective baselines to ensure fair and consistent comparisons.
Information Disentanglement. Once we obtain the feature representation $\mathbf{z}^{\mathrm{u}}$, we project it into two statistically independent components using two parallel $3\times 3$ convolutional layers with a channel ratio of $\alpha$. One is used to represent the content embedding, denoted as $\mathbf{z}_{c}^{\mathrm{u}}=f_{c}(\mathbf{z}^{\mathrm{u}})\in\mathbb{R}^{\alpha d}$, which provides view-agnostic scene structure. The other is employed for the viewpoint embedding, denoted as $\mathbf{z}_{v}^{\mathrm{u}}=f_{v}(\mathbf{z}^{\mathrm{u}})\in\mathbb{R}^{(1-\alpha)d}$, capturing view-specific attributes. To promote effective disentanglement, we introduce an intra-view independence constraint (Section 3.3) that minimizes the mutual information $\mathrm{MI}(Z_{c}^{\mathrm{u}};Z_{v}^{\mathrm{u}})$, where $Z^{\mathrm{u}}$ denotes the random variables of $\mathbf{z}^{\mathrm{u}}$, thereby encouraging statistical independence between content and viewpoint factors.
Cross-Reconstruction. While the independence constraint enhances factor separation, it may inadvertently lead to degenerate solutions or information loss. To mitigate this, we introduce an inter-view reconstruction constraint that encourages each factor to retain its intended information through cross-view image reconstruction. Specifically, we train two decoders, $D^{\mathrm{q}}$ and $D^{\mathrm{s}}$, to perform bidirectional reconstruction between paired views by swapping content and viewpoint embeddings, i.e., reconstructing each image using its own viewpoint and the content of the other. By enforcing accurate reconstruction from these hybrid embeddings, the model is incentivized to encode factor-specific information in each representation. This cross-view supervision not only prevents information collapse but also reinforces disentanglement. In the following, we describe the two specific constraints in CVD.
3.3 Intra-view Independence Constraint
To effectively disentangle content and viewpoint factors, we introduce an intra-view independence constraint that aims to minimize the statistical dependence between the two embeddings. Motivated by the principle that mutual information provides a fundamental measure of statistical dependence, we seek to encourage independent factorization by minimizing it between content and viewpoint.
Formally, the mutual information between $Z_{c}^{\mathrm{u}}$ and $Z_{v}^{\mathrm{u}}$ is defined as the Kullback-Leibler (KL) divergence between their joint distribution and the product of marginals:
$\mathrm{MI}(Z_{c}^{\mathrm{u}};Z_{v}^{\mathrm{u}})=\mathcal{D}_{\mathrm{KL}}\left(p(\mathbf{z}_{c}^{u},\mathbf{z}_{v}^{u})\;\middle\|\;p(\mathbf{z}_{c}^{u})\,p(\mathbf{z}_{v}^{u})\right).$ (2)
However, direct computation or optimization of mutual information is notoriously intractable in high-dimensional feature spaces due to the need for accurate estimation of joint and marginal densities. To circumvent this issue, we adopt the Sliced Wasserstein Distance (SWD) as a geometry-aware and sample-efficient proxy to promote independence. Specifically, we minimize the SWD between the empirical joint distribution $p(\mathbf{z}_{c}^{u},\mathbf{z}_{v}^{u})$ and the product of its marginals $p(\mathbf{z}_{c}^{u})\,p(\mathbf{z}_{v}^{u})$, which can be represented as:
$\mathcal{L}_{\text{iic}}^{\mathrm{u}}=\mathcal{SW}_{2}\left(p(\mathbf{z}_{c}^{u},\mathbf{z}_{v}^{u}),\;p(\mathbf{z}_{c}^{u})\otimes p(\mathbf{z}_{v}^{u})\right),$ (3)
where $\mathcal{SW}_{2}(\cdot,\cdot)$ denotes the Sliced Wasserstein-2 distance. We apply this constraint independently to both views, yielding $\mathcal{L}_{\text{iic}}^{\mathrm{q}}$ and $\mathcal{L}_{\text{iic}}^{\mathrm{s}}$. This constraint drives the separation of view-agnostic scene structure from view-specific attributes in a computationally tractable manner.
3.4 Inter-view Reconstruction Constraint
While the independence constraint promotes factor separation, it does not guarantee that each embedding retains the essential factor-specific information. In particular, relying solely on independence may lead to trivial solutions where either the content or viewpoint embedding becomes uninformative.
To address this, we introduce an inter-view reconstruction constraint that enforces information retention through cross-view image reconstruction. Specifically, we deploy decoders $D$ that reconstruct each image from a hybrid embedding composed of content from one view with the viewpoint from the other, which is denoted as:
^$\hat{\mathcal{I}}^{\mathrm{q}}=D^{\mathrm{q}}(\mathbf{z}_{c}^{\mathrm{s}},\mathbf{z}_{v}^{\mathrm{q}}),\quad\hat{\mathcal{I}}^{\mathrm{s}}=D^{\mathrm{s}}(\mathbf{z}_{c}^{\mathrm{q}},\mathbf{z}_{v}^{\mathrm{s}}).$ (4)
This constraint ensures that both $\mathbf{z}^{\mathrm{u}}_{c}$ and $\mathbf{z}^{\mathrm{u}}_{v}$ preserve distinct and sufficient information, thereby preventing representational collapse. Notably, the reconstruction is conditioned on the viewpoint, which governs spatial and geometric layout, while the content determines underlying scene structure. The reconstruction loss is defined as:
$\mathcal{L}^{\mathrm{u}}_{\text{irc}}=\|\mathcal{I}^{\mathrm{u}}-\hat{\mathcal{I}}^{\mathrm{u}}\|_{2}^{2}.$ (5)
where $\|\cdot\|_{2}^{2}$ denotes the mean squared error (MSE) between the original and reconstructed images. This loss is also applied to both views, resulting in $\mathcal{L}_{\text{irc}}^{\mathrm{q}}$ and $\mathcal{L}_{\text{irc}}^{\mathrm{s}}$, which ensure that the disentangled features preserve the necessary information to reconstruct their cross-view counterparts.
3.5 Training Objective
Following prior works [4, 11], we employ the standard InfoNCE loss for view-agnostic content consistency across views, denoted as $\mathcal{L}_{\text{loc}}$, which encourages content of matched query-satellite pairs to be close while pushing away mismatched pairs, and is defined as:
$\mathcal{L}_{\text{loc}}=-\log\frac{\exp(\mathbf{z}_{c}^{\mathrm{q}}\cdot\mathbf{z}_{c}^{\mathrm{s}}/\tau)}{\sum_{i=1}^{N}\exp(\mathbf{z}_{c}^{\mathrm{q}}\cdot\mathbf{z}_{c}^{\mathrm{i}}/\tau)},$ (6)
where $\tau$ is a temperature parameter that controls the sharpness of the similarity distribution. The overall training objective for CVD combines three losses: (1) an intra-view independence loss $\mathcal{L}^{\mathrm{u}}_{\text{iic}}$ promotes content and viewpoint independence, (2) an inter-view reconstruction loss $\mathcal{L}^{\mathrm{u}}_{\text{irc}}$ to ensure intended information preservation, and (3) a cross-view localization loss $\mathcal{L}_{\text{loc}}$ for discriminative alignment, which can be expressed as:
$\mathcal{L}_{\text{total}}=\lambda_{1}(\tfrac{1}{2}\mathcal{L}^{\mathrm{q}}_{\text{iic}}+\tfrac{1}{2}\mathcal{L}^{\mathrm{s}}_{\text{iic}})+\lambda_{2}(\tfrac{1}{2}\mathcal{L}^{\mathrm{q}}_{\text{irc}}+\tfrac{1}{2}\mathcal{L}^{\mathrm{s}}_{\text{irc}})+\lambda_{3}\mathcal{L}_{\text{loc}},$ (7)
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are hyperparameters that balance the three losses during training.
4 Experiments
4.1 Settings
Datasets. We evaluate our method on four representative CVGL benchmarks: University-1652 [28], SUES-200 [30], CVUSA [21], and CVACT_val [9], which collectively encompass various perspective combinations, including ground-view, drone-view, and satellite-view imagery. University-1652 comprises images from 1,652 university campuses captured in ground, drone, and satellite views. In our experiments, we adopt the drone-satellite setting, using 701 campuses for training, 701 for testing, along with 250 distractor samples. SUES-200 consists of real-world drone and satellite imagery from 200 scenes across four altitudes (150-300m), with 120 scenes for training and 80 for testing. CVUSA and CVACT_val provide large-scale ground-satellite image pairs, with 35,532 aligned in the training set for each. The former contains 8,884 test queries, while the latter offers a validation set of the same size.
Evaluation metrics. We adopt five standard retrieval metrics, including Average Precision (AP), Recall@K (K=1,5,10), and Recall@1%, to evaluate cross-view matching performance. Detailed definitions of metrics can be found in the Appendix.
Implementation details. The hyperparameters of the total training loss $\mathcal{L}_{\text{total}}$ are fixed across all models and datasets, with $\lambda_{1}=10$, $\lambda_{2}=0.2$, and $\lambda_{3}=1$. The temperature parameter $\tau$ is set to 0.05. To ensure fair comparisons, we strictly follow the original training configurations of each baseline, including optimizer type, learning rate schedule, batch size, etc., without any additional tuning. All experiments are conducted in PyTorch on an NVIDIA RTX 4090 GPU with 24 GB memory.
4.2 Results on University-1652
[…]
Table 1: Comparison between CVD and baseline methods on University-1652. $\dagger$’ represents that CVD is applied during training.
Method Image Size Backbone Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
AP R@1 R@5 R@10 R@1% AP R@1 R@5 R@10 R@1%
LPN 256$\times$256 ResNet50 77.26 73.87 88.84 92.58 93.01 73.55 85.28 89.27 91.15 98.29
[…]
SDPL 512$\times$512 ResNet50 86.45 84.13 94.36 96.45 96.72 82.17 89.44 92.58 93.58 99.29
[…]
MCCG 256$\times$256 ConvNeXt-T 90.63 88.92 96.44 97.63 97.77 88.73 93.15 95.72 96.72 99.57
[…]
Sample4Geo 384$\times$384 ConvNeXt-B 93.56 92.36 97.64 98.26 98.36 91.64 94.72 97.00 97.43 99.43
[…]
Game4Loc 384$\times$384 ViT-B 92.56 91.32 96.56 97.33 97.41 90.83 94.43 95.72 96.57 98.71
[…]
Table 2: Comparison between CVD and baseline methods on SUES-200. $\dagger$’ represents that CVD is applied during training.
Method Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
150m 200m 250m 300m 150m 200m 250m 300m
AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1
LPN 63.50 58.20 74.16 69.60 79.70 75.60 82.93 78.50 63.68 77.50 78.36 87.50 84.26 90.00 87.99 92.50
LPN $\dagger$ 64.24 59.77 74.39 70.38 81.14 77.50 84.26 79.38 64.88 78.40 80.13 88.74 85.50 93.75 89.72 93.41
FSRA 69.31 66.86 84.79 82.43 90.74 88.50 92.55 90.12 74.99 78.53 88.04 89.89 91.76 92.51 92.30 95.10
FSRA $\dagger$ 70.51 67.97 85.22 83.07 92.11 89.90 93.76 91.70 76.45 79.40 89.04 91.07 92.98 94.26 92.90 95.99
SDPL 76.64 72.07 84.98 81.92 89.53 87.05 92.34 90.35 70.28 80.00 80.57 86.25 85.64 88.75 87.43 90.00
SDPL $\dagger$ 77.11 75.30 85.19 83.00 91.25 88.99 92.97 90.86 71.43 80.80 81.84 87.10 85.81 91.25 88.52 91.25
MCCG 81.21 79.96 86.24 85.01 92.15 90.47 94.97 94.20 89.76 92.06 92.40 93.88 96.15 96.34 96.52 98.78
MCCG $\dagger$ 82.10 80.56 87.16 86.40 92.98 91.08 95.37 94.84 90.92 92.96 93.15 94.44 96.43 97.06 96.74 98.99
MBF 84.52 82.19 87.41 85.27 90.85 89.62 92.60 91.33 83.96 86.36 89.93 91.39 90.90 92.95 91.60 96.25
MBF $\dagger$ 85.24 83.19 88.20 85.52 91.90 90.46 93.48 92.09 85.11 86.97 91.00 92.51 92.23 93.80 91.80 96.98
Game4Loc 95.59 94.62 97.27 96.55 98.16 97.55 98.24 97.67 93.06 93.75 94.50 96.25 94.92 96.25 95.36 95.00
Game4Loc $\dagger$ 96.70 95.80 97.78 97.10 98.14 97.60 98.98 98.65 93.37 96.25 95.03 97.50 95.95 96.25 96.28 97.50
4.3 Results on SUES-200
[…]
4.4 Results on CVUSA and CVACT
We further evaluate CVD on CVUSA and CVACT, which involve ground-to-satellite matching under extreme viewpoint differences and substantial scene layout variation. As shown in Table 3, all baselines exhibit measurable improvements when trained with CVD, even though they already perform strongly. These improvements are particularly meaningful given the challenges posed by ground-view imagery, such as occlusion, illumination changes, and perspective distortion.
Table 3: Comparison between CVD and baseline methods on CVUSA and CVACT. $\dagger$’ represents that CVD is applied during training.
Method CVUSA CVACT_val
R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1%
LPN 85.43 95.20 96.91 99.40 78.86 89.97 92.07 95.34
LPN $\dagger$ 87.15 96.43 97.26 99.49 80.08 91.10 93.16 96.69
TransGeo 93.72 98.01 98.56 99.78 83.99 93.49 95.17 97.80
TransGeo $\dagger$ 94.35 98.83 99.07 99.80 84.96 94.61 95.90 98.44
GeoDTR 93.05 98.01 98.94 99.80 85.11 94.00 95.47 98.03
GeoDTR $\dagger$ 93.92 98.70 99.26 99.83 85.72 95.19 96.63 98.74
Sample4Geo 98.43 99.15 99.42 99.81 90.29 95.98 97.04 98.53
Sample4Geo $\dagger$ 98.67 99.40 99.78 99.89 90.94 96.80 97.51 98.81
By explicitly separating view-specific conflicting information, CVD enhances the consistency of content representations across modalities. For example, Sample4Geo improves by +0.24% and +0.65% in R@1 on CVUSA and CVACT, respectively, while GeoDTR achieves +0.77% and +0.61% gains in R@10.
Figure 3: Visualization of disentangled content and viewpoint attention maps produced by CVD.
4.5 Visualization of Disentangled Embeddings
Figure 3 visualizes attention maps derived from the disentangled embeddings produced by CVD. Viewpoint embeddings predominantly focus on peripheral, view-specific elements such as vegetation, shadows, and roads, which are often inconsistent across modalities. In contrast, content embeddings attend to semantically stable regions like building structures, providing reliable and discriminative cues for cross-view matching. This separation facilitates CVD to learn robust, viewpoint-agnostic content representations, thereby improving retrieval under significant viewpoint disparity. More visualizations are given in the Appendix.
Table 4: Ablation of the components of CVD.
Exp. Methods (ResNet50) D $\rightarrow$ S S $\rightarrow$ D
AP R@1 AP R@1
1 Baseline LPN 77.26 73.87 73.55 85.28
2 w/o two constraints 74.62 71.83 69.85 83.49
3 Exp.1 + iic loss 77.95 73.95 73.97 85.24
4 Exp.1 + irc loss 78.28 74.62 74.63 85.45
5 CVD (LPN) 78.99 75.78 74.89 85.88
6 Baseline Game4Loc 72.90 68.49 66.53 83.16
7 w/o two constraints 69.71 65.85 64.51 81.19
8 Exp.6 + iic loss 73.29 69.19 66.70 83.20
9 Exp.7 + irc loss 73.60 69.78 67.03 83.61
10 CVD (Game4Loc) 74.31 70.46 67.60 83.67
Table 5: Ablation of content-viewpoint ratio.
Split Ratio D $\rightarrow$ S S $\rightarrow$ D
AP R@1 AP R@1
$\alpha=1/3$ 78.13 74.90 74.08 84.31
$\alpha=1/2$ 78.99 75.78 74.89 85.88
$\alpha=3/4$ 78.15 74.76 73.28 86.59
No squeeze 78.06 74.85 74.48 84.85
Table 6: Influence of different $\tau$ in InfoNCE.
Method D $\rightarrow$ S S $\rightarrow$ D
AP R@1 AP R@1
$\tau=0.07$ 78.89 75.74 74.28 85.03
$\tau=0.05$ 78.99 75.78 74.89 85.88
$\tau=0.03$ 78.98 75.76 74.61 85.43
Bi-InfoNCE 78.72 75.14 74.36 85.21
4.6 Ablation Study
Effectiveness of CVD. We conduct ablation studies on University-1652 to evaluate the contribution of each component in CVD, as summarized in Section 4.5. To ensure fair comparison, we adopt two representative pipelines, LPN and Game4Loc, which both use a shared ResNet50 backbone. Removing both constraints (Exp.2 and 7) leads to a notable performance drop, while individually adding the intra-view independence (Exp.3 and 8) or inter-view reconstruction constraint (Exp.4 and 9) yields consistent gains. The best results are obtained when both constraints are jointly applied (Exp.5 and 10), confirming that explicitly factorizing content and viewpoint is essential for robust cross-view alignment.
[…]
Temperature Parameter Analysis. As shown in Section 4.5, model performance remains stable across a range of temperature values $\tau$. This insensitivity indicates that the performance gains primarily result from effective representation disentanglement rather than contrastive loss tuning.
Table 7: Generalization ablation when trained on University-1652 and tested on SUES-200. $\dagger$’ represents that CVD is applied during training.
Method Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
150m 200m 250m 300m 150m 200m 250m 300m
AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1
LPN 42.83 36.70 52.99 46.72 59.42 53.62 62.15 56.55 25.30 30.00 34.36 38.75 38.53 42.50 43.92 53.75
LPN $\dagger$ 44.93 38.58 56.68 50.50 62.16 55.73 64.89 58.60 29.21 32.50 39.18 42.50 45.98 52.50 52.52 60.00
Game4Loc 82.39 78.85 88.57 86.10 90.31 88.17 90.94 88.75 75.29 80.00 81.31 88.75 84.31 88.75 86.40 92.50
Game4Loc $\dagger$ 86.12 82.87 91.77 89.70 93.59 91.92 94.44 92.92 75.37 87.50 84.25 90.00 87.97 95.00 90.21 95.00
4.7 Generalisation Capabilities
To evaluate the generalization ability of CVD to unseen scenes, we train models on University-1652 and directly evaluate them on SUES-200 without any fine-tuning. As shown in Table 7, CVD significantly improves performance across all metrics for both LPN (ResNet-50) and Game4Loc (ViT-B). For example, it boosts R@1 by up to +6.25% for LPN and +7.5% for Game4Loc, even without access to the target domain during training. Remarkably, these cross-dataset improvements exceed those typically obtained via in-domain training on SUES-200, which yields only improvements of 1-3%. These results highlight that disentangling content and viewpoint leads to more transferable representations that generalize effectively across regions and views.
Figure 4: Cross-view reconstructions with matched (left column) and unrelated (right column) viewpoints.
4.8 Effectiveness of Content-Viewpoint Disentanglement
To further evaluate the effectiveness of disentangling content and viewpoint, we conduct a cross-view recombination experiment using the learned embeddings. As illustrated in Figure 4, each row shows two reconstructed heatmaps, which share the same content but differ in the viewpoint. In the top row, the same satellite content embedding is combined with either a matched (left) or unrelated (right) drone viewpoint; in the bottom row, drone content is paired with matched and unrelated satellite viewpoints, respectively. Despite the change in viewpoint, the resulting attention maps exhibit similar attention patterns, highlight semantically consistent regions. These results provide qualitative evidence that CVD effectively disentangles the two factors: content embeddings encode view-agnostic structural information, while viewpoint embeddings modulate geometric and contextual variations. More visualizations are provided in the Appendix.
5 Conclusion and Limitations
In this paper, we revisit cross-view geo-localization (CVGL) from a manifold learning perspective and propose CVD, a unified framework that explicitly disentangles content and viewpoint within the visual representation. CVD adopts an embed-disentangle-reconstruct paradigm, guided by intra-view independence and inter-view reconstruction constraints, to promote content- and viewpoint-specific encoding. Extensive experiments demonstrate that CVD consistently enhances both localization accuracy and generalization across diverse CVGL pipelines. These findings underscore the value of explicitly separating content and viewpoint in CVGL, and offer insights toward developing more robust and generalizable CVGL systems. A discussion of limitations and potential future directions is provided in the Appendix.

### v2

@@
-Robust Cross-View Geo-Localization via Content-Viewpoint Disentanglement
+Robust Drone-View Geo-Localization via Content-Viewpoint Disentanglement
 Abstract
-Cross-view geo-localization (CVGL) aims to match images of the same geographic location captured from different perspectives, such as drones and satellites.
-Despite recent advances, CVGL remains highly challenging due to significant appearance changes and spatial distortions caused by viewpoint variations.
-Existing methods typically assume that cross-view images can be directly aligned within a shared feature space by maximizing feature similarity through contrastive learning.
+Drone-view geo-localization (DVGL) aims to match images of the same geographic location captured from drone and satellite perspectives.
+Despite recent advances, DVGL remains challenging due to significant appearance changes and spatial distortions caused by viewpoint variations.
+Existing methods typically assume that drone and satellite images can be directly aligned in a shared feature space via contrastive learning.
 Nonetheless, this assumption overlooks the inherent conflicts induced by viewpoint discrepancies, resulting in extracted features containing inconsistent information that hinders precise localization.
-In this study, we take a manifold learning perspective and model the feature space of cross-view images as a composite manifold jointly governed by content and viewpoint information.
-Building upon this insight, we propose CVD, a new CVGL framework that explicitly disentangles content and viewpoint factors.
-To promote effective disentanglement, we introduce two constraints: (i) An intra-view independence constraint, which encourages statistical independence between the two factors by minimizing their mutual information.
-(ii) An inter-view reconstruction constraint that reconstructs each view by cross-combining content and viewpoint from paired images, ensuring factor-specific semantics are preserved.
-As a plug-and-play module, CVD can be seamlessly integrated into existing geo-localization pipelines.
-Extensive experiments on four benchmarks, i.e., University-1652, SUES-200, CVUSA, and CVACT, demonstrate that CVD consistently improves both localization accuracy and generalization across multiple baselines.
-11footnotetext: Corresponding author: wangdi@xidian.edu.cn
+In this study, we take a manifold learning perspective and model the feature space of cross-view images as a composite manifold jointly governed by content and viewpoint.
+Building upon this insight, we propose CVD, a new DVGL framework that explicitly disentangles content and viewpoint factors.
+To promote effective disentanglement, we introduce two constraints: (i) an intra-view independence constraint that encourages statistical independence between the two factors by minimizing their mutual information; and (ii) an inter-view reconstruction constraint that reconstructs each view by cross-combining content and viewpoint from paired images, ensuring factor-specific semantics are preserved.
+As a plug-and-play module, CVD integrates seamlessly into existing DVGL pipelines and reduces inference latency.
+Extensive experiments on University-1652 and SUES-200 show that CVD exhibits strong robustness and generalization across various scenarios, viewpoints and altitudes, with further evaluations on CVUSA and CVACT confirming consistent improvements.
 1 Introduction
-The widespread deployment of autonomous vehicles, drones, and other intelligent systems has placed growing demands on navigation and localization technologies.
-To ensure safe and reliable operation, high-precision positioning services have become a fundamental requirement.
-Cross-view geo-localization (CVGL), an onboard technique independent of external communication infrastructure, offers a promising solution by estimating absolute geospatial coordinates in the absence of conventional localization signals (e.g., GPS) [18, 3, 25, 8].
-Given a query image, typically captured from a ground or drone perspective, the goal is to find a matching satellite image from a georeferenced database to infer the query’s location.
-Most existing approaches formulate CVGL as an image retrieval task [17, 23], training deep neural networks (DNNs) to learn visual similarity across different views.
-However, the viewpoint disparity between query and satellite imagery introduces severe spatial distortions and appearance variations, making robust matching inherently challenging.
 Figure 1: Comparison between previous methods and our CVD.
@@
 Right: Our method learns disentangled representations by mapping inputs onto two submanifolds corresponding to content $\mathcal{M}_{c}$ and viewpoint $\mathcal{M}_{v}$.
-This separation is enforced via two constraints (see Section 3.3 and Section 3.4), promoting effective disentanglement and thereby enhancing cross-view matching performance.
-Recent efforts aim to alleviate the viewpoint discrepancy between query and satellite images.
-A common strategy is to apply predefined geometric transformations to satellite images, such as polar transformation or orthorectification, aligning their spatial layout with that of the query view.
-However, the effectiveness of these methods relies on prior knowledge of the geometric relationship between views, and may degrade when the query is not spatially centered within the satellite image [31].
-Despite recent advances in CVGL, many existing methods [9, 22, 31, 4, 27] still follow a common training paradigm, i.e., directly applying contrastive learning to pull the features of positive pairs closer and push those of negative pairs away.
+This separation is enforced via two constraints (see Sec.
+3.3 and Sec.
+3.4), promoting effective disentanglement and thereby enhancing cross-view matching performance.
+The widespread deployment of drones and other intelligent systems has placed growing demands on navigation and localization technologies.
+To ensure safe and reliable operation, high-precision positioning services have become a fundamental requirement.
+Drone-view geo-localization (DVGL), an onboard technique independent of external communication infrastructure, offers a promising solution by estimating absolute geospatial coordinates in the absence of conventional localization signals (e.g., GPS) [sun2023cross, dai2023vision, zhang2024aligning, lin2015learning].
+Given a drone image, the goal is to find a matching satellite image from a georeferenced database to infer the drone’s location.
+Most existing approaches formulate DVGL as an image retrieval task [sun2021multisensor, ye2024cross2], training deep neural networks (DNNs) to learn visual similarity across different views.
+However, the viewpoint disparity between drone and satellite imagery introduces severe spatial distortions and appearance variations, making robust matching inherently challenging.
+Recent efforts aim to alleviate the viewpoint discrepancy between drone and satellite images.
+A common strategy is to employ predefined geometric transformations to satellite images, such as polar transformation or orthorectification, aligning their spatial layout with that of the drone view.
+However, the effectiveness of these methods relies on prior knowledge of the geometric relationship between the two views and may degrade when the drone image is not spatially centered within the satellite image [zhu2022transgeo].
+Despite recent advances in DVGL, many existing methods [liu2019lending, yang2021cross, zhu2022transgeo, deuser2023sample4geo, zhang2023cross] still follow a common training paradigm, i.e., directly applying contrastive learning to pull the features of positive pairs closer and push those of negative pairs away.
 While various architectural or training refinements have been proposed, these methods largely overlook the semantic inconsistencies introduced by drastic viewpoint differences.
-These inconsistencies disrupt the alignment between positive pairs and ultimately limit the cross-view matching performance.
-In this paper, we revisit the CVGL task from a manifold learning perspective by modeling the feature space of cross-view images as a composite manifold jointly governed by content and viewpoint factors.
-As illustrated in Figure 1 (left), prior methods can be interpreted as learning representations on a composite manifold $\mathcal{M}$, where both factors are inherently entangled.
-Such entanglement introduces viewpoint-induced conflicts into the learned representations, thereby undermining the robustness of contrastive alignment.
-To address this limitation, we propose CVD (Content-Viewpoint Disentanglement), a generic CVGL framework that explicitly factorizes the feature space into two submanifolds: a content submanifold and a viewpoint submanifold (see Figure 1 (right)).
-The content encodes view-agnostic structural information, while the viewpoint captures view-specific appearance variations.
-As shown in Figure 2 (Training Stage), CVD adopts an embed-disentangle-reconstruct paradigm.
-Each image is first embedded into a shared feature space, then projected onto independent content and viewpoint submanifolds, and finally recombined in an image reconstruction task.
-To facilitate effective disentanglement, we impose two dedicated constraints: an intra-view independence constraint and an inter-view reconstruction constraint.
-The former encourages statistical independence between content and viewpoint by minimizing their mutual information, while the latter ensures that each factor retains its intended semantics by reconstructing images using the content of one view and the viewpoint of the other.
-In addition, we apply a standard contrastive loss (e.g., InfoNCE) to align content representations of matched pairs across views.
-To the best of our knowledge, CVD is the first CVGL framework to explicitly disentangle content and viewpoint information.
-Unlike existing methods that rely on architectural redesign, CVD serves as a plug-and-play module that can be seamlessly integrated into existing pipelines.
-It introduces no additional cost during inference while substantially improving matching accuracy and robustness.
+These inconsistencies disrupt the alignment between positive pairs and ultimately limit the drone-satellite matching performance.
+In this paper, we revisit the DVGL task from a manifold learning perspective by modeling the feature space of cross-view images as a composite manifold jointly governed by content and viewpoint factors.
+As illustrated in Fig.
+1 (left), prior methods can be interpreted as learning representations on a single composite manifold $\mathcal{M}$ where both factors are entangled.
+Such entanglement injects viewpoint-induced conflicts in the learned representations, undermining the robustness of contrastive alignment.
+To address this limitation, we propose CVD (Content-Viewpoint Disentanglement), a general DVGL framework that explicitly factorizes the feature space into two submanifolds: content $\mathcal{M}_{c}$ and viewpoint $\mathcal{M}_{v}$ (see Fig.
+1 (right)).
+The content encodes view-agnostic geo-structural information, while the viewpoint captures view-specific appearance variations.
+As shown in Fig.
+2, CVD adopts an embed-disentangle-reconstruct paradigm: each image is first embedded into a shared feature space, then projected onto independent content and viewpoint submanifolds, and finally recombined via an image reconstruction task.
+To facilitate the disentanglement, we impose two dedicated constraints: an intra-view independence constraint and an inter-view reconstruction constraint.
+The former encourages statistical independence between content and viewpoint by minimizing their mutual information, while the latter preserves the intended semantics of each factor by reconstructing one view using the content from the drone (or satellite) image and the viewpoint from its satellite (or drone) counterpart.
+In addition, we apply standard contrastive loss (e.g., InfoNCE) to align content representations of matched drone-satellite pairs.
+To the best of our knowledge, CVD is the first DVGL framework to explicitly disentangle content and viewpoint.
+Unlike approaches that require bespoke architectural redesigns, CVD integrates as a plug-and-play module into existing pipelines and lowering inference-time overhead.
+Extensive experiments on four benchmarks, i.e.
+University-1652, SUES-200, CVUSA and CVACT, demonstrate consistent gains over multiple baselines, improving robustness to viewpoint and altitude changes and generalization under scene shifts.
 In summary, our contributions are as follows:
-  • We revisit the cross-view geo-localization (CVGL) task from a manifold learning perspective and propose CVD, the first framework that explicitly disentangles content and viewpoint to suppress viewpoint-induced conflicting information.
-  • We introduce two constraints to facilitate effective disentanglement: an intra-view independence constraint that facilitates the independence between content and viewpoint, and an inter-view reconstruction constraint that ensures each factor retains its intended semantics.
-  • CVD can be seamlessly integrated into existing pipelines, incurs no additional inference overhead, and consistently improves cross-view matching performance and generalization across multiple challenging benchmarks.
+  • We revisit the DVGL task from a manifold-learning perspective and propose CVD, the first framework that explicitly disentangles content and viewpoint to suppress viewpoint-induced conflicting information.
+  • We introduce two constraints to facilitate disentanglement: an intra-view independence constraint that facilitates the independence between content and viewpoint, and an inter-view reconstruction constraint that ensures each factor preserves its intended semantics
+  • CVD integrates seamlessly into existing pipelines, shortens inference time, and consistently improves cross-view matching performance, enabling efficient deployment in practice.
+  • Extensive experiments show that CVD improves the robustness and generalization of DVGL pipelines across diverse scenarios, viewpoints, and altitudes.
 2 Related Work
-CVGL has witnessed significant progress with the availability of large-scale geo-tagged datasets and advances in deep learning.
+2.1 Visual-based Geo-Localization
+Visual-based geo-localization (VGL) has witnessed significant progress with the availability of large-scale geo-tagged datasets and advances in deep learning.
 Existing methods predominantly follow a Siamese-based framework and can be broadly categorized into three research directions: data augmentation strategies, architectural innovations, and feature representation learning.
-2.1 Data Augmentation Strategies
-Data augmentation has become a widely adopted strategy in CVGL tasks [19, 12, 26, 6].
-To address cross-view misalignment, Liu et al.[9] incorporate camera orientation as auxiliary input, while SAFA[15] applies a polar transformation to align aerial and ground-level panoramas.
-Several methods further improve robustness by diversifying input image appearances.
-For instance, Lv et al.[10] simulate extreme weather conditions by injecting synthetic variations into University-1652.
+Data Augmentation Strategies.
+Data augmentation has become a widely adopted strategy in VGL tasks [vo2016localizing, rodrigues2021these, zhang2024benchmarking, li2024unleashing].
+To address cross-view misalignment, Liu et al.[liu2019lending] incorporate camera orientation as auxiliary input, while SAFA [shi2019spatial] applies a polar transformation to align aerial and ground-level panoramas.
+CVGlobal [ye2024cross] introduces a panoramic BEV transformation based on the ground-plane assumption and geometric constraints, effectively reducing the gap between street panoramas and satellite imagery.
+Similarly, Video2BEV [ju2025video2bev] transforms drone videos into BEV representations, facilitating better alignment with satellite imagery.
 More recently, training-aware data sampling has emerged as a complementary strategy.
-Sample4Geo[4] introduces two curriculum-driven strategies: one leveraging geographically adjacent samples for easier early-stage alignment, and another mining hard negatives to refine the decision boundary.
-Game4Loc [5] proposes a mutual-exclusion sampling mechanism that enforces strict decorrelation between positive and negative pairs, thereby enhancing contrastive supervision in cross-view matching.
-2.2 Backbone Innovations
-In parallel, a line of work focuses on designing more powerful visual backbones to enhance localization performance [13, 29].
-For example, L2LTR [22] exploits self-attention to model long-range dependencies, effectively reducing visual ambiguity in cross-view.
-RK-Net [7] introduces a lightweight unit-difference attention module that enables joint learning of dense features and salient keypoints, without requiring additional annotations.
-SAIG [33] proposes an efficient backbone tailored for CVGL by replacing the MLP blocks in standard Transformers with spatially-aware mixing layers and low-dimensional projections, yielding a more compact and structured representation.
-2.3 Feature Representation Learning
-Many studies focus on learning more effective visual representations to enhance cross-view matching [32, 11, 30, 24].
+Sample4Geo [deuser2023sample4geo] introduces two curriculum-driven strategies: one leveraging geographically adjacent samples for easier early-stage alignment, and another mining hard negatives to refine the decision boundary.
+DenseUAV [dai2023vision] integrates metric learning with mutual supervision, effectively reducing modality discrepancy and improving feature discriminability.
+Game4Loc [ji2025game4loc] proposes a mutual-exclusion sampling mechanism that enforces strict decorrelation between positive and negative pairs, thereby enhancing contrastive supervision in cross-view matching.
+Backbone Innovations.
+In parallel, a line of work focuses on designing more powerful visual backbones to enhance localization performance [rodrigues2022global, zhu2023modern].
+For example, L2LTR [yang2021cross] exploits self-attention to model long-range dependencies, effectively reducing visual ambiguity in cross-view.
+RK-Net [lin2022joint] introduces a lightweight unit-difference attention module that enables joint learning of dense features and salient keypoints, without requiring additional annotations.
+SAIG [zhu2023simple] proposes an efficient backbone tailored for VGL by replacing the MLP blocks in standard Transformers with spatially-aware mixing layers and low-dimensional projections, yielding a more compact and structured representation.
+Feature Representation Learning.
+Many studies focus on learning more effective visual representations to enhance cross-view matching [zhu2021vigor, mi2024congeo, zhu2023sues, zhai2017predicting, zeng2022geo, ti2024tirsaa].
 A common strategy involves refining alignment mechanisms between views.
-For instance, Shi et al.[16] proposed a dynamic similarity matching network to estimate directional alignment, thereby reducing cross-view discrepancies.
-FSRA[2] leverages transformer-based heatmaps to perform region-level alignment, while LPN [20] incorporates contextual cues via a square-ring partitioning strategy to improve part-based representations.
-SDPL [1] builds upon LPN by introducing a shifting-fusion mechanism to generate multiple complementary part sets, which are then adaptively aggregated to enhance robustness against spatial shifts and scale variations.
+For instance, Shi et al.[shi2020looking] proposed a dynamic similarity matching network to estimate directional alignment, thereby reducing cross-view discrepancies.
+FSRA [dai2021transformer] leverages transformer-based heatmaps to perform region-level alignment, while LPN [wang2021each] incorporates contextual cues via a square-ring partitioning strategy to improve part-based representations.
+SDPL [chen2024sdpl] builds upon LPN by introducing a shifting-fusion mechanism to generate multiple complementary part sets, which are then adaptively aggregated to enhance robustness against spatial shifts and scale variations.
 Several methods also incorporate inductive priors modeling to enhance feature expressiveness.
-TransGeo [31] adopts a non-uniform cropping strategy that discards low-information regions while reallocating resolution to semantically salient areas, enhancing accuracy without increasing computational cost.
-GeoDTR [27] incorporates spatial layout priors through a geometry-aware extraction module, combined with counterfactual training to exploit structural information.
-MCCG [14] enriches feature diversity by jointly modeling spatial and channel-wise attentions.
+FRGeo [zhang2024aligning] enhances cross-view alignment by explicitly recombining spatial features to reduce geometric ambiguities.
+TransGeo [zhu2022transgeo] adopts a non-uniform cropping strategy that discards low-information regions while reallocating resolution to semantically salient areas, enhancing accuracy without increasing computational cost.
+MCCG [shen2023mccg] enriches feature diversity by jointly modeling spatial and channel-wise attentions.
 Although prior methods perform well across various scenarios, many rely on auxiliary components that increase inference-time overhead.
-In contrast, we propose a plug-and-play training paradigm that explicitly disentangles content and viewpoint from raw features, enhancing cross-view correspondence without introducing additional parameters or latency at inference.
-Figure 2: Overview of the proposed CVD.
-3 Methodology
+In contrast, we propose a plug-and-play training paradigm that explicitly disentangles content and viewpoint from raw feature representations, thereby effectively enhancing cross-view correspondence and reducing inference latency and computational cost.
+2.2 Disentangled Representation Learning
+Disentangled representation learning (DRL) aims to learn representations that identify and disentangle the underlying factors hidden in observable data [wang2024disentangled].
+Owing to the resulting interpretability, controllability, and robustness, it has seen broad adoption in computer vision [li2025fd2, ruan2022adaptive, wang2016multimodal, chen2024disendreamer, cheng2024disentangled], natural language processing [zhao2024disentangled, bao2019generating, cheng2020improving], recommender systems [ma2019learning], and graph learning [wang2020disentangled, zhang2023dyted, wang2022disencite], with gains on many downstream tasks.
+For example, Zou et al.
+[zou2020joint] tackle cross-domain person re-identification by jointly disentangling ID-related and ID-unrelated subspaces and restricting adaptation to the former, thereby improving transferability.
+DisCo [Wang_2024_CVPR] introduces a disentangled-control architecture that separates subject, background, and pose, enabling compositional and generalizable dance video synthesis.
+Wang et al.
+[wang2024towards] present a frequency-domain disentanglement framework for UAV object detection that employs two learnable filters to isolate domain-invariant from domain-specific spectra, leading to stronger domain generalization.
+Recent studies have begun to incorporate DRL into VGL tasks.
+GeoDTR [zhang2023cross] introduces a geometry-aware layout extractor to separate geometric cues from raw appearance features, thereby improving cross-view localization.
+However, it leaves unaddressed the viewpoint-induced conflicts persisting in both appearance and layout representations, thereby hindering cross-view correspondence.
+In this work, we explicitly disentangle content from viewpoint, and employ independence and reconstruction constraints to suppress such conflicts, yielding a cleaner, view-agnostic content representation.
+3 Methods
 3.1 Problem Formulation
-Considering a set of image pairs $\{(\mathcal{I}_{i}^{\mathrm{q}},\mathcal{I}_{i}^{\mathrm{s}})\}_{i=1}^{N}$, where superscripts $\mathrm{q}$ and $\mathrm{s}$ denote query (e.g., drone or ground) and satellite images, respectively, and $N$ is the number of pairs.
+Considering a set of image pairs $\{(\mathcal{I}_{i}^{\mathrm{d}},\mathcal{I}_{i}^{\mathrm{s}})\}_{i=1}^{N}$, where superscripts $\mathrm{d}$ and $\mathrm{s}$ denote drone and satellite images, respectively, and $N$ is the number of pairs.
 Each pair depicts the same geographic location.
-In the CVGL task, given a query image $\mathcal{I}_{q}^{\mathrm{q}}$ with index $q$, the objective is to retrieve its best-matching satellite image $\mathcal{I}_{b}^{\mathrm{s}}$ from the georeferenced database, where $b\in\{1,\ldots,N\}$.
+In the DVGL task, given a drone image $\mathcal{I}_{d}^{\mathrm{d}}$ with index $d$, the objective is to retrieve its best-matching satellite image $\mathcal{I}_{b}^{\mathrm{s}}$ from the georeferenced database, where $b\in\{1,\ldots,N\}$.
 Most existing methods rely on learning a representation function $f(\cdot)$ that embeds images from different viewpoints into a shared feature space, allowing matching pairs to be identified through feature distance minimization.
@@
 These are composed via a function $f:\mathcal{M}_{c}\times\mathcal{M}_{v}\rightarrow\mathcal{Z}$ that maps both factors to a point in the feature space.
-Assuming statistical independence between the factors, i.e., $p(C,V)=p(C)\,p(V)$, the resulting distribution over $\mathcal{Z}$ is given by the push-forward measure $f_{\#}(p(C)\times p(V))$.
-This assumption is well-aligned with the CVGL task, where identical scenes may be observed under diverse perspectives.
+Assuming statistical independence between the factors, i.e., p(C,V)=p(C)p(V)p(C,V)=p(C)\,p(V), the resulting distribution over $\mathcal{Z}$ is given by the push-forward measure $f_{\#}(p(C)\times p(V))$.
+This assumption is well-aligned with the DVGL task, where identical scenes may be observed under diverse perspectives.
 To realize the above formulation, we design manifold encoders that decompose each image $\mathcal{I}^{\mathrm{u}}$ into two distinct representations: a content embedding $f_{c}(\mathcal{I}^{\mathrm{u}})\in\mathcal{M}_{c}$ and a viewpoint embedding $f_{v}(\mathcal{I}^{\mathrm{u}})\in\mathcal{M}_{v}$.
-Cross-view matching is subsequently performed in the $\mathcal{M}_{c}$ by retrieving the nearest neighbor of a query image $\mathcal{I}_{q}^{\mathrm{q}}$ via its content embedding:
-$b=\underset{i\in\{1,\ldots,N\}}{\arg\min}d(f_{c}(\mathcal{I}_{q}^{\mathrm{q}}),f_{c}(\mathcal{I}_{i}^{\mathrm{s}})),$ (1)
+Cross-view matching is subsequently performed in the $\mathcal{M}_{c}$ by retrieving the nearest neighbor of a drone image $\mathcal{I}_{q}^{\mathrm{d}}$ via its content embedding:
+$b=\underset{i\in\{1,\ldots,N\}}{\arg\min}d(f_{c}(\mathcal{I}_{d}^{\mathrm{d}}),f_{c}(\mathcal{I}_{i}^{\mathrm{s}})),$ (1)
 where $d(\cdot,\cdot)$ denotes a distance metric.
-For notation compactness, we will use superscript $u$  [^2] We adopt this convention throughout the paper.
-for cases that apply to both query ($\mathrm{q}$) and satellite ($\mathrm{s}$) views.
+For notation compactness, we will use superscript uu [^2] We adopt this convention throughout the paper.
+for cases that apply to both drone ($\mathrm{d}$) and satellite ($\mathrm{s}$) views.
 3.2 Proposed Methodology
-As illustrated in Figure 2, CVD adopts a Siamese architecture consisting of two symmetric branches for the query and satellite views.
+As illustrated in Fig.
+2, CVD adopts a Siamese architecture consisting of two symmetric branches for the drone and satellite views.
 Each branch comprises three sequential components: manifold embedding, information disentanglement, and cross-reconstruction.
 Manifold Embedding.
-Each input image $\mathcal{I}^{\mathrm{u}}$ ($u\in\{\mathrm{q},\mathrm{s}\}$) is first processed by manifold encoders $E$, yielding a raw $d$-dimensional feature representations $\mathbf{z}^{\mathrm{u}}=E(\mathcal{I}^{\mathrm{u}})\in\mathbb{R}^{d}$.
+Each input image $\mathcal{I}^{\mathrm{u}}$ ($u\in\{\mathrm{d},\mathrm{s}\}$) is first processed by manifold encoders $E$, yielding a raw $d$-dimensional feature representations $\mathbf{z}^{\mathrm{u}}=E(\mathcal{I}^{\mathrm{u}})\in\mathbb{R}^{d}$.
 Owing to the nature of DNN encoders, the distribution of $\mathbf{z}^{\mathrm{u}}$ can be viewed as residing on a composite manifold jointly governed by content and viewpoint information.
 Since our primary focus is on the training paradigm, we adopt the same DNN encoders as those used in the respective baselines to ensure fair and consistent comparisons.
+Figure 2: Overview of the proposed CVD.
 Information Disentanglement.
@@
 The other is employed for the viewpoint embedding, denoted as $\mathbf{z}_{v}^{\mathrm{u}}=f_{v}(\mathbf{z}^{\mathrm{u}})\in\mathbb{R}^{(1-\alpha)d}$, capturing view-specific attributes.
-To promote effective disentanglement, we introduce an intra-view independence constraint (Section 3.3) that minimizes the mutual information $\mathrm{MI}(Z_{c}^{\mathrm{u}};Z_{v}^{\mathrm{u}})$, where $Z^{\mathrm{u}}$ denotes the random variables of $\mathbf{z}^{\mathrm{u}}$, thereby encouraging statistical independence between content and viewpoint factors.
+To promote effective disentanglement, we introduce an intra-view independence constraint (Sec.
+3.3) that minimizes the mutual information $\mathrm{MI}(Z_{c}^{\mathrm{u}};Z_{v}^{\mathrm{u}})$, where $Z^{\mathrm{u}}$ denotes the random variables of $\mathbf{z}^{\mathrm{u}}$, thereby encouraging statistical independence between content and viewpoint factors.
 Cross-Reconstruction.
 While the independence constraint enhances factor separation, it may inadvertently lead to degenerate solutions or information loss.
-To mitigate this, we introduce an inter-view reconstruction constraint that encourages each factor to retain its intended information through cross-view image reconstruction.
-Specifically, we train two decoders, $D^{\mathrm{q}}$ and $D^{\mathrm{s}}$, to perform bidirectional reconstruction between paired views by swapping content and viewpoint embeddings, i.e., reconstructing each image using its own viewpoint and the content of the other.
+To mitigate this, we introduce an inter-view reconstruction constraint (Sec.
+3.4) that encourages each factor to retain its intended information through cross-view image reconstruction.
+Specifically, we train two decoders, $D^{\mathrm{d}}$ and $D^{\mathrm{s}}$, to perform bidirectional reconstruction between paired views by swapping content and viewpoint embeddings, i.e., reconstructing each image using its own viewpoint and the content of the other.
 By enforcing accurate reconstruction from these hybrid embeddings, the model is incentivized to encode factor-specific information in each representation.
@@
 where $\mathcal{SW}_{2}(\cdot,\cdot)$ denotes the Sliced Wasserstein-2 distance.
-We apply this constraint independently to both views, yielding $\mathcal{L}_{\text{iic}}^{\mathrm{q}}$ and $\mathcal{L}_{\text{iic}}^{\mathrm{s}}$.
+We apply this constraint independently to both views, yielding $\mathcal{L}_{\text{iic}}^{\mathrm{d}}$ and $\mathcal{L}_{\text{iic}}^{\mathrm{s}}$.
 This constraint drives the separation of view-agnostic scene structure from view-specific attributes in a computationally tractable manner.
@@
 While the independence constraint promotes factor separation, it does not guarantee that each embedding retains the essential factor-specific information.
-In particular, relying solely on independence may lead to trivial solutions where either the content or viewpoint embedding becomes uninformative.
+In particular, relying solely on independence may lead to trivial solutions where either the content or the viewpoint embedding becomes uninformative.
 To address this, we introduce an inter-view reconstruction constraint that enforces information retention through cross-view image reconstruction.
 Specifically, we deploy decoders $D$ that reconstruct each image from a hybrid embedding composed of content from one view with the viewpoint from the other, which is denoted as:
-^$\hat{\mathcal{I}}^{\mathrm{q}}=D^{\mathrm{q}}(\mathbf{z}_{c}^{\mathrm{s}},\mathbf{z}_{v}^{\mathrm{q}}),\quad\hat{\mathcal{I}}^{\mathrm{s}}=D^{\mathrm{s}}(\mathbf{z}_{c}^{\mathrm{q}},\mathbf{z}_{v}^{\mathrm{s}}).$ (4)
+$\hat{\mathcal{I}}^{\mathrm{d}}=D^{\mathrm{d}}(\mathbf{z}_{c}^{\mathrm{s}},\mathbf{z}_{v}^{\mathrm{d}}),\quad\hat{\mathcal{I}}^{\mathrm{s}}=D^{\mathrm{s}}(\mathbf{z}_{c}^{\mathrm{d}},\mathbf{z}_{v}^{\mathrm{s}}).$ (4)
 This constraint ensures that both $\mathbf{z}^{\mathrm{u}}_{c}$ and $\mathbf{z}^{\mathrm{u}}_{v}$ preserve distinct and sufficient information, thereby preventing representational collapse.
@@
 where $\|\cdot\|_{2}^{2}$ denotes the mean squared error (MSE) between the original and reconstructed images.
-This loss is also applied to both views, resulting in $\mathcal{L}_{\text{irc}}^{\mathrm{q}}$ and $\mathcal{L}_{\text{irc}}^{\mathrm{s}}$, which ensure that the disentangled features preserve the necessary information to reconstruct their cross-view counterparts.
+This loss is also applied to both views, resulting in $\mathcal{L}_{\text{irc}}^{\mathrm{d}}$ and $\mathcal{L}_{\text{irc}}^{\mathrm{s}}$, which ensure that the disentangled features preserve the necessary information to reconstruct their cross-view counterparts.
 3.5 Training Objective
-Following prior works [4, 11], we employ the standard InfoNCE loss for view-agnostic content consistency across views, denoted as $\mathcal{L}_{\text{loc}}$, which encourages content of matched query-satellite pairs to be close while pushing away mismatched pairs, and is defined as:
-$\mathcal{L}_{\text{loc}}=-\log\frac{\exp(\mathbf{z}_{c}^{\mathrm{q}}\cdot\mathbf{z}_{c}^{\mathrm{s}}/\tau)}{\sum_{i=1}^{N}\exp(\mathbf{z}_{c}^{\mathrm{q}}\cdot\mathbf{z}_{c}^{\mathrm{i}}/\tau)},$ (6)
+Following prior works [deuser2023sample4geo, mi2024congeo], we employ the standard InfoNCE loss for view-agnostic content consistency across views, denoted as $\mathcal{L}_{\text{loc}}$, which encourages content of matched drone-satellite pairs to be close while pushing away mismatched pairs, and is defined as:
+$\mathcal{L}_{\text{loc}}=-\log\frac{\exp(\mathbf{z}_{c}^{\mathrm{d}}\cdot\mathbf{z}_{c}^{\mathrm{s}}/\tau)}{\sum_{i=1}^{N}\exp(\mathbf{z}_{c}^{\mathrm{d}}\cdot\mathbf{z}_{c}^{\mathrm{i}}/\tau)},$ (6)
 where $\tau$ is a temperature parameter that controls the sharpness of the similarity distribution.
 The overall training objective for CVD combines three losses: (1) an intra-view independence loss $\mathcal{L}^{\mathrm{u}}_{\text{iic}}$ promotes content and viewpoint independence, (2) an inter-view reconstruction loss $\mathcal{L}^{\mathrm{u}}_{\text{irc}}$ to ensure intended information preservation, and (3) a cross-view localization loss $\mathcal{L}_{\text{loc}}$ for discriminative alignment, which can be expressed as:
-$\mathcal{L}_{\text{total}}=\lambda_{1}(\tfrac{1}{2}\mathcal{L}^{\mathrm{q}}_{\text{iic}}+\tfrac{1}{2}\mathcal{L}^{\mathrm{s}}_{\text{iic}})+\lambda_{2}(\tfrac{1}{2}\mathcal{L}^{\mathrm{q}}_{\text{irc}}+\tfrac{1}{2}\mathcal{L}^{\mathrm{s}}_{\text{irc}})+\lambda_{3}\mathcal{L}_{\text{loc}},$ (7)
-where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are hyperparameters that balance the three losses during training.
+$\mathcal{L}_{\text{total}}=\lambda_{1}(\tfrac{1}{2}\mathcal{L}^{\mathrm{d}}_{\text{iic}}+\tfrac{1}{2}\mathcal{L}^{\mathrm{s}}_{\text{iic}})+\lambda_{2}(\tfrac{1}{2}\mathcal{L}^{\mathrm{d}}_{\text{irc}}+\tfrac{1}{2}\mathcal{L}^{\mathrm{s}}_{\text{irc}})+\mathcal{L}_{\text{loc}},$ (7)
+where $\lambda_{1}$ and $\lambda_{2}$ are two loss-balancing hyperparameters.
 4 Experiments
@@
 Datasets.
-We evaluate our method on four representative CVGL benchmarks: University-1652 [28], SUES-200 [30], CVUSA [21], and CVACT_val [9], which collectively encompass various perspective combinations, including ground-view, drone-view, and satellite-view imagery.
-University-1652 comprises images from 1,652 university campuses captured in ground, drone, and satellite views.
+We evaluate our method on four representative CVGL benchmarks: University-1652 [zheng2020university] comprises images from 1,652 university campuses captured in ground, drone, and satellite views.
 In our experiments, we adopt the drone-satellite setting, using 701 campuses for training, 701 for testing, along with 250 distractor samples.
-SUES-200 consists of real-world drone and satellite imagery from 200 scenes across four altitudes (150-300m), with 120 scenes for training and 80 for testing.
-CVUSA and CVACT_val provide large-scale ground-satellite image pairs, with 35,532 aligned in the training set for each.
-The former contains 8,884 test queries, while the latter offers a validation set of the same size.
+SUES-200 [zhu2023sues] consists of real-world drone and satellite imagery from 200 scenes across four altitudes (150-300m), with 120 scenes for training and 80 for testing.
+Together, University-1652 and SUES-200 span diverse scene types, flight altitudes, and viewing directions, providing a rigorous testbed to assess the robustness and generalization of our approach.
+To further validate effectiveness, we additionally evaluate on the CVUSA [workman2015wide] and CVACT_val [liu2019lending] datasets (ground $\rightarrow$ satellite), which provide 35,532 aligned training pairs each; the former contains 8,884 test queries, and the latter offers a test set of the same size.
 Evaluation metrics.
-We adopt five standard retrieval metrics, including Average Precision (AP), Recall@K (K=1,5,10), and Recall@1%, to evaluate cross-view matching performance.
-Detailed definitions of metrics can be found in the Appendix.
+We adopt five retrieval metrics, including Average Precision (AP), Recall@K (K=1,5,10), and Recall@1%, to evaluate cross-view matching performance.
+Definitions of these metrics are provided in the Appendix.
+Baselines.
+To evaluate the effectiveness of our method, we compare it against several SOTA methods, including LPN [wang2021each], FSRA [dai2021transformer], TransGeo [zhu2022transgeo], MCCG [shen2023mccg], Sample4Geo [deuser2023sample4geo], SDPL [chen2024sdpl], Game4Loc [ji2025game4loc] and GeoDTR [zhang2023cross].
 Implementation details.
-The hyperparameters of the total training loss $\mathcal{L}_{\text{total}}$ are fixed across all models and datasets, with $\lambda_{1}=10$, $\lambda_{2}=0.2$, and $\lambda_{3}=1$.
+The hyperparameters of the total training loss are fixed across all experiments, with $\lambda_{1}=10$, and $\lambda_{2}=0.2$.
 The temperature parameter $\tau$ is set to 0.05.
-To ensure fair comparisons, we strictly follow the original training configurations of each baseline, including optimizer type, learning rate schedule, batch size, etc., without any additional tuning.
-All experiments are conducted in PyTorch on an NVIDIA RTX 4090 GPU with 24 GB memory.
-4.2 Results on University-1652
-[…]
-Table 1: Comparison between CVD and baseline methods on University-1652. $\dagger$’ represents that CVD is applied during training.
+To ensure fair comparisons, we strictly follow the original training configurations of each baseline, including optimizer type, learning rate schedule, etc., without any additional tuning.
+All experiments are conducted in PyTorch on an NVIDIA 4090 GPU.
+Each experiment is repeated three times with different random seeds, and the mean results are reported to ensure statistical reliability.
+Table 1: Comparison of baselines and their CVD-enhanced counterparts (marked with $\dagger$) on the University-1652 dataset.
 Method Image Size Backbone Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
 AP R@1 R@5 R@10 R@1% AP R@1 R@5 R@10 R@1%
-LPN 256$\times$256 ResNet50 77.26 73.87 88.84 92.58 93.01 73.55 85.28 89.27 91.15 98.29
+LPN $\times256$ ResNet50 77.26 73.87 88.84 92.58 93.01 73.55 85.28 89.27 91.15 98.29
 […]
-SDPL 512$\times$512 ResNet50 86.45 84.13 94.36 96.45 96.72 82.17 89.44 92.58 93.58 99.29
+FSRA $\times256$ ViT-S 84.24 81.62 93.06 95.19 95.43 80.99 87.87 90.87 92.87 98.43
 […]
-MCCG 256$\times$256 ConvNeXt-T 90.63 88.92 96.44 97.63 97.77 88.73 93.15 95.72 96.72 99.57
+SDPL $\times512$ ResNet50 86.45 84.13 94.36 96.45 96.72 82.17 89.44 92.58 93.58 99.29
 […]
-Sample4Geo 384$\times$384 ConvNeXt-B 93.56 92.36 97.64 98.26 98.36 91.64 94.72 97.00 97.43 99.43
+MCCG $\times256$ ConvNeXt-T 90.63 88.92 96.44 97.63 97.77 88.73 93.15 95.72 96.72 99.57
 […]
-Game4Loc 384$\times$384 ViT-B 92.56 91.32 96.56 97.33 97.41 90.83 94.43 95.72 96.57 98.71
+Sample4Geo $\times384$ ConvNeXt-B 93.56 92.36 97.64 98.26 98.36 91.64 94.72 97.00 97.43 99.43
 […]
-Table 2: Comparison between CVD and baseline methods on SUES-200. $\dagger$’ represents that CVD is applied during training.
+Game4Loc $\times384$ ViT-B 92.56 91.32 96.56 97.33 97.41 90.83 94.43 95.72 96.57 98.71
+[…]
+Table 2: Comparison of baselines and their CVD-enhanced counterparts (marked with $\dagger$) on the SUES-200 dataset.
 Method Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
@@
 LPN 63.50 58.20 74.16 69.60 79.70 75.60 82.93 78.50 63.68 77.50 78.36 87.50 84.26 90.00 87.99 92.50
-LPN $\dagger$ 64.24 59.77 74.39 70.38 81.14 77.50 84.26 79.38 64.88 78.40 80.13 88.74 85.50 93.75 89.72 93.41
-FSRA 69.31 66.86 84.79 82.43 90.74 88.50 92.55 90.12 74.99 78.53 88.04 89.89 91.76 92.51 92.30 95.10
-FSRA $\dagger$ 70.51 67.97 85.22 83.07 92.11 89.90 93.76 91.70 76.45 79.40 89.04 91.07 92.98 94.26 92.90 95.99
+[…]
+FSRA 82.69 78.70 88.66 85.65 91.26 88.95 93.40 91.50 83.65 93.75 90.01 93.75 91.67 97.50 92.49 95.00
+[…]
 SDPL 76.64 72.07 84.98 81.92 89.53 87.05 92.34 90.35 70.28 80.00 80.57 86.25 85.64 88.75 87.43 90.00
-SDPL $\dagger$ 77.11 75.30 85.19 83.00 91.25 88.99 92.97 90.86 71.43 80.80 81.84 87.10 85.81 91.25 88.52 91.25
+[…]
 MCCG 81.21 79.96 86.24 85.01 92.15 90.47 94.97 94.20 89.76 92.06 92.40 93.88 96.15 96.34 96.52 98.78
 MCCG $\dagger$ 82.10 80.56 87.16 86.40 92.98 91.08 95.37 94.84 90.92 92.96 93.15 94.44 96.43 97.06 96.74 98.99
-MBF 84.52 82.19 87.41 85.27 90.85 89.62 92.60 91.33 83.96 86.36 89.93 91.39 90.90 92.95 91.60 96.25
-MBF $\dagger$ 85.24 83.19 88.20 85.52 91.90 90.46 93.48 92.09 85.11 86.97 91.00 92.51 92.23 93.80 91.80 96.98
+Sample4Geo 96.08 94.75 97.69 96.75 98.38 97.25 98.41 97.20 95.60 96.25 96.41 96.25 96.54 96.25 96.57 97.50
+[…]
 Game4Loc 95.59 94.62 97.27 96.55 98.16 97.55 98.24 97.67 93.06 93.75 94.50 96.25 94.92 96.25 95.36 95.00
-Game4Loc $\dagger$ 96.70 95.80 97.78 97.10 98.14 97.60 98.98 98.65 93.37 96.25 95.03 97.50 95.95 96.25 96.28 97.50
-4.3 Results on SUES-200
 […]
-4.4 Results on CVUSA and CVACT
-We further evaluate CVD on CVUSA and CVACT, which involve ground-to-satellite matching under extreme viewpoint differences and substantial scene layout variation.
-As shown in Table 3, all baselines exhibit measurable improvements when trained with CVD, even though they already perform strongly.
-These improvements are particularly meaningful given the challenges posed by ground-view imagery, such as occlusion, illumination changes, and perspective distortion.
-Table 3: Comparison between CVD and baseline methods on CVUSA and CVACT. $\dagger$’ represents that CVD is applied during training.
+4.2 Main Results
+[…]
+[…]
+Table 3: Cross-dataset generalization results comparing baselines and their CVD-enhanced counterparts (marked with $\dagger$), trained on the University-1652 dataset and directly tested on the SUES-200 dataset.
+Method Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
+150m 200m 250m 300m 150m 200m 250m 300m
+AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1
+LPN 42.83 36.70 52.99 46.72 59.42 53.62 62.15 56.55 25.30 30.00 34.36 38.75 38.53 42.50 43.92 53.75
+LPN $\dagger$ 44.93 38.58 56.68 50.50 62.16 55.73 64.89 58.60 29.21 32.50 39.18 42.50 45.98 52.50 52.52 60.00
+FSRA 58.22 52.45 67.10 61.87 70.64 66.07 71.99 67.50 50.95 58.75 59.07 66.25 61.07 62.50 61.98 63.75
+[…]
+SDPL 38.52 32.80 47.30 41.33 52.38 46.72 53.62 48.50 25.26 27.50 35.05 37.50 41.74 46.25 43.85 48.75
+[…]
+MCCG 74.99 70.85 86.04 83.20 90.84 88.90 93.38 91.85 59.85 63.75 74.65 81.25 79.87 83.75 81.17 86.25
+[…]
+Sample4Geo 64.24 57.62 74.45 69.00 81.22 77.02 85.48 81.90 83.85 88.75 90.09 92.50 91.68 96.25 93.51 95.00
+[…]
+Game4Loc 82.39 78.85 88.57 86.10 90.31 88.17 90.94 88.75 75.29 80.00 81.31 88.75 84.31 88.75 86.40 92.50
+[…]
+Table 4: Comparison of baselines and their CVD-enhanced counterparts (marked with $\dagger$) on the CVUSA and CVACT datasets.
 Method CVUSA CVACT_val
@@
 LPN 85.43 95.20 96.91 99.40 78.86 89.97 92.07 95.34
-LPN $\dagger$ 87.15 96.43 97.26 99.49 80.08 91.10 93.16 96.69
+[…]
 TransGeo 93.72 98.01 98.56 99.78 83.99 93.49 95.17 97.80
-TransGeo $\dagger$ 94.35 98.83 99.07 99.80 84.96 94.61 95.90 98.44
+[…]
 GeoDTR 93.05 98.01 98.94 99.80 85.11 94.00 95.47 98.03
-GeoDTR $\dagger$ 93.92 98.70 99.26 99.83 85.72 95.19 96.63 98.74
+[…]
+Game4Loc 98.12 99.08 99.46 99.83 90.19 96.00 97.12 98.60
+[…]
 Sample4Geo 98.43 99.15 99.42 99.81 90.29 95.98 97.04 98.53
-Sample4Geo $\dagger$ 98.67 99.40 99.78 99.89 90.94 96.80 97.51 98.81
+[…]
+Results on University-1652 $\rightarrow$ SUES-200.
+To evaluate the generalization ability of CVD to unseen scenes, we train models on University-1652 and directly evaluate them on SUES-200 without any fine-tuning.
+As shown in Tab.
+3, CVD significantly improves performance across all metrics for both LPN (ResNet-50) and Game4Loc (ViT-B).
+For example, it boosts R@1 by up to +6.25% for LPN and +7.5% for Game4Loc, even without access to the target domain during training.
+Remarkably, these cross-dataset improvements exceed those typically obtained via in-domain training on SUES-200, which yields only improvements of 1-3%.
+These results highlight that disentangling content and viewpoint leads to more transferable representations that generalize effectively across regions and views.
+Results on CVUSA and CVACT.
+To further examine the generality of CVD beyond the drone-view setting, we also evaluate it on CVUSA and CVACT, which involve ground-to-satellite matching under extreme viewpoint differences and substantial scene layout variation.
+As shown in Tab.
+4, all baselines exhibit continuous improvements when trained with CVD, even though they already perform strongly.
+These improvements are particularly meaningful given the challenges posed by ground-view imagery, such as occlusion, illumination changes, and perspective distortion.
 By explicitly separating view-specific conflicting information, CVD enhances the consistency of content representations across modalities.
 For example, Sample4Geo improves by +0.24% and +0.65% in R@1 on CVUSA and CVACT, respectively, while GeoDTR achieves +0.77% and +0.61% gains in R@10.
-Figure 3: Visualization of disentangled content and viewpoint attention maps produced by CVD.
-4.5 Visualization of Disentangled Embeddings
-Figure 3 visualizes attention maps derived from the disentangled embeddings produced by CVD.
-Viewpoint embeddings predominantly focus on peripheral, view-specific elements such as vegetation, shadows, and roads, which are often inconsistent across modalities.
-In contrast, content embeddings attend to semantically stable regions like building structures, providing reliable and discriminative cues for cross-view matching.
-This separation facilitates CVD to learn robust, viewpoint-agnostic content representations, thereby improving retrieval under significant viewpoint disparity.
-More visualizations are given in the Appendix.
-Table 4: Ablation of the components of CVD.
-Exp. Methods (ResNet50) D $\rightarrow$ S S $\rightarrow$ D
+Qualitative Results.
+Appendix Sec.
+10.1, we provide qualitative comparisons of cross-view retrieval results on University-1652 using both CNN-based LPN and Transformer-based Game4Loc.
+The results intuitively confirm that CVD disentangles view-agnostic content from viewpoint-specific variations, enabling more reliable cross-view correspondence.
+[…]
+Figure 3: Qualitative results of cross-view reconstruction on the University-1652 and SUES-200 datasets.
+Table 5: Ablation studies on: (a) Effect of CVD’s components. (b) Content-viewpoint split ratio $\alpha$. (c) Analysis of different $\tau$ in InfoNCE.
+Exp. Methods (ResNet50) Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
 AP R@1 AP R@1
@@
 2 w/o two constraints 74.62 71.83 69.85 83.49
-3 Exp.1 + iic loss 77.95 73.95 73.97 85.24
-4 Exp.1 + irc loss 78.28 74.62 74.63 85.45
+3 Exp.1 + $\mathcal{L}_{\text{iic}}$ 77.95 73.95 73.97 85.24
+4 Exp.1 + $\mathcal{L}_{\text{irc}}$ 78.28 74.62 74.63 85.45
 5 CVD (LPN) 78.99 75.78 74.89 85.88
@@
 7 w/o two constraints 69.71 65.85 64.51 81.19
-8 Exp.6 + iic loss 73.29 69.19 66.70 83.20
-9 Exp.7 + irc loss 73.60 69.78 67.03 83.61
+8 Exp.6 + $\mathcal{L}_{\text{iic}}$ 73.29 69.19 66.70 83.20
+9 Exp.7 + $\mathcal{L}_{\text{irc}}$ 73.60 69.78 67.03 83.61
 10 CVD (Game4Loc) 74.31 70.46 67.60 83.67
-Table 5: Ablation of content-viewpoint ratio.
-Split Ratio D $\rightarrow$ S S $\rightarrow$ D
+Split Ratio Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
 AP R@1 AP R@1
@@
 No squeeze 78.06 74.85 74.48 84.85
-Table 6: Influence of different $\tau$ in InfoNCE.
-Method D $\rightarrow$ S S $\rightarrow$ D
+Method Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
 AP R@1 AP R@1
@@
 Bi-InfoNCE 78.72 75.14 74.36 85.21
+Table 6: Effect of different reconstruction loss functions on the University-1652 dataset.
+[…]
+AP R@1 R@5 R@10 R@1% AP R@1 R@5 R@10 R@1%
+MSE 19.38 0.4608 93.95 92.94 97.59 98.28 98.32 91.92 94.86 96.71 97.14 99.57
+SSIM 19.21 0.4601 93.19 92.05 97.26 97.87 97.90 91.78 94.55 96.40 96.19 99.56
+Perceptual 19.06 0.4589 92.70 91.45 96.82 97.84 97.91 90.85 94.14 97.42 97.26 99.29
+4.3 Effectiveness of Disentangled Strategy
+To evaluate our disentanglement strategy, we conduct cross-view image reconstruction (Drone $\rightarrow$ Satellite) on University-1652 and SUES-200.
+As illustrated in Fig.
+3, the visualization results show that the reconstructed outputs consistently preserve the global layout, semantic topology, and structural relations of the original scenes, despite the loss of certain high-frequency and color details.
+For example, the upper-right sample faithfully recovers the circular central plaza and the relative arrangement of surrounding buildings.
+This indicates that CVD successfully learned meaningful content and viewpoint representations, thereby verifying the effectiveness of disentangling.
+Note. Additional visualizations are provided in Appendix Secs. 10.2 to 10.4, including more cross-view reconstructions and attention maps.
 4.4 Ablation Study
-Effectiveness of CVD.
-We conduct ablation studies on University-1652 to evaluate the contribution of each component in CVD, as summarized in Section 4.5.
+Effect of CVD’s Components.
+We conduct ablation studies on the University-1652 dataset to evaluate the contribution of each component in CVD, as summarized in Tab.
+5(a).
 To ensure fair comparison, we adopt two representative pipelines, LPN and Game4Loc, which both use a shared ResNet50 backbone.
@@
 […]
-Temperature Parameter Analysis.
-As shown in Section 4.5, model performance remains stable across a range of temperature values $\tau$.
+Analysis of Temperature Parameters.
+As illustrated in Tab.
+5(c), model performance remains stable across a range of temperature values $\tau$.
 This insensitivity indicates that the performance gains primarily result from effective representation disentanglement rather than contrastive loss tuning.
-Table 7: Generalization ablation when trained on University-1652 and tested on SUES-200. $\dagger$’ represents that CVD is applied during training.
-Method Drone $\rightarrow$ Satellite Satellite $\rightarrow$ Drone
-150m 200m 250m 300m 150m 200m 250m 300m
-AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1
-LPN 42.83 36.70 52.99 46.72 59.42 53.62 62.15 56.55 25.30 30.00 34.36 38.75 38.53 42.50 43.92 53.75
-LPN $\dagger$ 44.93 38.58 56.68 50.50 62.16 55.73 64.89 58.60 29.21 32.50 39.18 42.50 45.98 52.50 52.52 60.00
-Game4Loc 82.39 78.85 88.57 86.10 90.31 88.17 90.94 88.75 75.29 80.00 81.31 88.75 84.31 88.75 86.40 92.50
-Game4Loc $\dagger$ 86.12 82.87 91.77 89.70 93.59 91.92 94.44 92.92 75.37 87.50 84.25 90.00 87.97 95.00 90.21 95.00
-4.7 Generalisation Capabilities
-To evaluate the generalization ability of CVD to unseen scenes, we train models on University-1652 and directly evaluate them on SUES-200 without any fine-tuning.
-As shown in Table 7, CVD significantly improves performance across all metrics for both LPN (ResNet-50) and Game4Loc (ViT-B).
-For example, it boosts R@1 by up to +6.25% for LPN and +7.5% for Game4Loc, even without access to the target domain during training.
-Remarkably, these cross-dataset improvements exceed those typically obtained via in-domain training on SUES-200, which yields only improvements of 1-3%.
-These results highlight that disentangling content and viewpoint leads to more transferable representations that generalize effectively across regions and views.
-Figure 4: Cross-view reconstructions with matched (left column) and unrelated (right column) viewpoints.
-4.8 Effectiveness of Content-Viewpoint Disentanglement
-To further evaluate the effectiveness of disentangling content and viewpoint, we conduct a cross-view recombination experiment using the learned embeddings.
-As illustrated in Figure 4, each row shows two reconstructed heatmaps, which share the same content but differ in the viewpoint.
-In the top row, the same satellite content embedding is combined with either a matched (left) or unrelated (right) drone viewpoint; in the bottom row, drone content is paired with matched and unrelated satellite viewpoints, respectively.
-Despite the change in viewpoint, the resulting attention maps exhibit similar attention patterns, highlight semantically consistent regions.
-These results provide qualitative evidence that CVD effectively disentangles the two factors: content embeddings encode view-agnostic structural information, while viewpoint embeddings modulate geometric and contextual variations.
-More visualizations are provided in the Appendix.
-5 Conclusion and Limitations
-In this paper, we revisit cross-view geo-localization (CVGL) from a manifold learning perspective and propose CVD, a unified framework that explicitly disentangles content and viewpoint within the visual representation.
-CVD adopts an embed-disentangle-reconstruct paradigm, guided by intra-view independence and inter-view reconstruction constraints, to promote content- and viewpoint-specific encoding.
-Extensive experiments demonstrate that CVD consistently enhances both localization accuracy and generalization across diverse CVGL pipelines.
-These findings underscore the value of explicitly separating content and viewpoint in CVGL, and offer insights toward developing more robust and generalizable CVGL systems.
-A discussion of limitations and potential future directions is provided in the Appendix.
+Effect of Reconstruction Losses.
+We compare MSE, SSIM, and perceptual losses.
+As shown in Tab.
+6, MSE achieves the highest PSNR/SSIM and best retrieval accuracy, indicating that pixel-level fidelity better preserves view consistency.
+By contrast, SSIM and perceptual losses relax strict photometric fidelity and show higher tolerance to small misalignments, which may reduce geometric consistency and slightly affect retrieval performance.
+We therefore adopt MSE in all experiments.
+Note.
+We provide additional ablations in the Appendix, including robustness to training data scale (Tab.
+8), the number of SWD projections (Tab.
+9), and the effect of loss-balancing weights (Tab.
+11), among others.
+5 Conclusion
+In this paper, we revisit drone-view geo-localization (DVGL) from a manifold-learning perspective and propose CVD, a unified framework that explicitly disentangles content and viewpoint within visual representations.
+CVD follows an embed-disentangle-reconstruct paradigm, guided by intra-view independence and inter-view reconstruction constraints, to promote content- and viewpoint-specific encoding.
+Extensive experiments show that CVD consistently improves cross-view matching accuracy across diverse pipelines and enhances robustness and generalization under varied scenarios, viewpoints, and altitudes, while achieving lower inference latency.
+These results underscore the value of separating content and viewpoint for DVGL and point toward more robust and generalizable DVGL systems.
+We discuss limitations and directions for future work in the Appendix.